How much should you invest in a continuously compounded account at an annually interest rate of 6% if you want exactly $8000 after four years?
Answer by lwsshak3(11628)

How much should you invest in a continuously compounded account at an annually interest rate of 6% if you want exactly $8000 after four years?
***
Formula for a continuous compounding account: A=Pe^rt, P=initial investment, r=interest rate, t=years, A=amt after t-yrs
For given problem:
r=.06
t=4
A=8000
..
P=A/e^rt=8000/e^(.08*4)=8000/e^(.32)=5809.19
How much should you invest in the continuously compounded account? $5809